A truck off the from rest and cover a distance of 120m on a straight road in 20 second

Calculate the acceleration

1
Expert's answer
2021-02-22T11:08:24-0500

Total distance covered(s) = 120 m

Time taken(t) = 20 s

Initial velocity of particle(u) = 0 m/s

Final velocity(v) = "\dfrac{s}{t}=\dfrac{120}{20}m/s"

= "6m\/s"

Acceleration = "\dfrac{v-u}{t}=\dfrac{6}{20}m/s^2"

= "0.3m\/s^2"